OpenStudy (anonymous):

The width of a rectangle is fixed at 6 cm. Determine those length for which the area will be less than 138cm2

OpenStudy (anonymous):

Width is 3 feet. Length if 5 feet. Area = L*W = 3*5 = 15 square feet.